pt 3 (1 point) The expression can be written as eft), where f(x) is a function of . Find f(x). f(x) = x+2
Answer #1

We know

\large \frac{e^m}{e^n}=e^{m-n}

so

\large \frac{e^{x-3}}{e^{x-8}}=e^{(x-3)-(x-8)}=e^{(x-3-x+8)}=e^5

so

f(x)= 5

so answer is 5